The volume of a sphere that is 1 inch 2 feet in diameter is __ cubic feet?
Answer by checkley77(12844)

V=4/3(PIR^3)
V=4/3(3.14*12.5^3) (NOTE 1 IN+2FT=25 IN. 25/2=12.5 IS THE RADIUS)
V=4/3(3.14*1,953.125
V=4/3*6,132.8125
V=8,177.0833 IN^3 ANS. OR 4.7321 FT^3